Differentiate the following and calculate the value of π’…π’š at the value of x stated: 𝒅𝐱

1. π’š=πŸπ’™πŸ‘ +πŸ’π’™πŸ βˆ’πŸπ’™+πŸ• at 𝒙=βˆ’πŸ

2. π’š=πŸ‘π’™πŸ’ βˆ’πŸ“π’™πŸ‘ +πŸ’π’™πŸ βˆ’π’™+πŸ’ at 𝒙=πŸ‘


1
Expert's answer
2021-05-07T09:00:21-0400
  1. "𝒚=𝟐𝒙^𝟑 +𝟒𝒙^𝟐 −𝟐𝒙+𝟕\\"

Differentiate this equation with respect to x


"\dfrac{dy}{dx}=6x^2+8x-2"


So "\dfrac{dy}{dx}" at "x=-2" so put the value of "x=-2" in above equation


"\Rightarrow 6\times(-2)^2+8(-2)-2=6"


2."\ud835\udc9a=\ud835\udfd1\ud835\udc99^\ud835\udfd2 \u2212\ud835\udfd3\ud835\udc99^\ud835\udfd1 +\ud835\udfd2\ud835\udc99^\ud835\udfd0 \u2212\ud835\udc99+\ud835\udfd2"


Differentiate this equation with respect to x


"\dfrac{dy}{dx}=12x^3-15x^2+8x-1"


So "\dfrac{dy}{dx}" at "x=3" so put the value of "x=3" in above equation


"\Rightarrow 12\times(3)^3-15\times(3)^2+8\times(3)-1" "=212"